A certain element exists in a face-centered cubic structure. The atomic radius of an atom of this element is 121 pm. Calculate the density of this element in g/cm3. (Assume the element has a molar mass of 59.3 g/mol.)
Answer #1

FCC

a = 2\sqrt{}2 r

r   = 121pm = 121*10^-10 cm

a    = 2\sqrt{}2 *121*10^-10 cm

       = 342.2*10^-10 cm

Z     = 4 for FCC

M     = 59.3g/mole

d     = ZM/Na^3

        = 4*59.3/6.023*10^23 *(342.2*10^-10)^3

        = 9.83g/cm^3 >>>>answer